OpenStudy (anonymous):

y > 2x^2+4x + 6 how do i provide 2 ordered pairs which satisfy the quadratic inequality?? Please Help!!

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

do you know how?

jimthompson5910 (jim_thompson5910):

if you plug in something like `x = 0` into `y = 2x^2 + 4x + 6`, what value of y do you get?

OpenStudy (anonymous):

i was goin to say thet

OpenStudy (anonymous):

there is the answer

OpenStudy (anonymous):

It Has to be less than zero

OpenStudy (anonymous):

which is why im confused?

OpenStudy (anonymous):

Ordered pair solutions to equations - Khan Academy Video for y > 2x^2+4x + 6 how do i provide 2 ordered pairs which satisfy the quadratic inequality?

OpenStudy (anonymous):

look at the vid it will help

OpenStudy (anonymous):

Okay Hold On

OpenStudy (anonymous):

This is inequalities not equations

OpenStudy (anonymous):

oh sorry

jimthompson5910 (jim_thompson5910):

Sorry I got distracted Replace every x with 0 y = 2x^2 + 4x + 6 y = 2*0^2 + 4*0 + 6 y = 6 so the ordered pair (0,6) is on the graph of y = 2x^2 + 4x + 6 what you can do is simply bump up the y coordinate to get to (0,7). This point is guaranteed to be in the solution region because the shaded solution region is above the parabolic curve Repeat these steps for other x values like x = 1, x = 2, etc